# DIP: Unsupervised Dense In-Context Post-training of Visual Representations

### v2

@@ 3.1 Dense In-Context Training
During training, we sample an in-context task $\mathcal{T}$ from dataset $\mathcal{D}$. Each task $\mathcal{T}$ consists of $K$ support examples $\mathcal{S}=\{(X_{s_{i}},Y_{s_{i}})\}_{i=1}^{K}$ and one query example $(X_{q},Y_{q})$, where $X_{q}$ and $X_{s_{i}}$ are images, and $Y_{q}$ and $Y_{s_{i}}$ are their one-hot semantic segmentation pseudo-labels. Among the $K$ support examples, one is a “positive” example sharing pseudo-classes with the query, while the remaining $K-1$ are “distractor” examples unlikely to share any pseudo-class. Since our method is unsupervised, these tasks and pseudo-labels are automatically constructed (see Sec. 3.2).
@@ 3.2 Automatic Dense In-Context Task Construction
Selecting examples for in-context tasks.  To generate in-context tasks, we pair each query image with one “positive” support image and $K$-1 “distractor” support images.
[…]
For the $K$-1 “distractor” examples, we randomly sample images from the current mini-batch during training. Including distractors encourages the model to distinguish relevant from irrelevant content, improving feature discriminability.
@@ 4.1 Experimental Setup
$\frac{1}{128}$$\frac{1}{64}$$\frac{1}{8}$$\frac{1}{1}$606060656565707070757575808080Fraction of DatamIoU (%)(a) PascalVOC $\frac{1}{128}$$\frac{1}{64}$$\frac{1}{8}$$\frac{1}{1}$262626282828303030323232343434363636383838404040424242Fraction of Data(b) ADE20K
@@ 4.2 Retrieval-based Scene Understanding
Post-training other base models.  In Tab. 4, we evaluate in-context segmentation on ADE20K and PascalVOC using CLIP [61] and MAE [31] as base models, before and after DIP post-training. Our method consistently improves performance across all tested models, showing its versatility. MAE improves by +33.4+33.4+ 33.4 mIoU on PascalVOC. These gains transform MAE from weak to improved at in-context segmentation, highlighting our approach’s effectiveness in enhancing dense features.
@@ 5 Conclusion
We introduced DIP, a novel unsupervised post-training method that enhances dense image representations for in-context scene understanding. Leveraging meta-learning principles and automatically generating in-context tasks with pseudo-segmentations using Stable Diffusion, our approach avoids complex self-distillation architectures. DIP is computationally efficient ($<$9h on an A100) and generalizes well across downstream dense retrieval tasks, including semantic segmentation and depth prediction. It outperforms both the initial pretrained vision encoder and prior state-of-the-art post-training methods, providing a simpler and more effective solution for improving dense representations in vision models.
@@ Appendix A Implementation details
During post-training, we fine-tune only the last three transformer blocks of the pretrained ViT encoder $f(\cdot)$ while keeping the remaining layers frozen. Our MLP projector $h(\cdot)$ consists of two linear layers with a non-linear activation function GELU [35]. Hidden feature dimension is set to $7D$ and fixed output dimension 6144. $\ell_{2}$-normalization is applied at the output of the MLP. We set the temperature parameter $\tau$ of the softmax operator to 0.070.070.07 (see Tab. 7). Our ablation study demonstrates that the method is robust to the choice of temperature, with values ranging from 0.030.030.03 to 0.090.090.09 yielding similar performance on the PascalVOC dataset. During post-training our support set consists of 1 positive and 7 “distractor” examples (8 total examples). During training, we use the AdamW [49] optimizer with a learning rate of $2.25\times 10^{-7}$ and a weight decay of 0.050.050.05. We train for 555 epochs. We employ cosine learning rate schedule.
@@ Appendix B Additional quantitative results
$10^{5}$$10^{6}$$10^{7}$343434363636383838404040424242Memory SizemIoU (%)(a) ViT-S/14 $10^{5}$$10^{6}$$10^{7}$343434363636383838404040424242Memory Size(a) ViT-B/14
@@ 1 $\frac{1}{8}$ $\frac{1}{64}$ $\frac{1}{128}$
111 $\frac{1}{8}$ $\frac{1}{64}$ $\frac{1}{128}$ 111 $\frac{1}{8}$ $\frac{1}{64}$ $\frac{1}{128}$

### v3

@@ 3.1 Dense In-Context Training
During training, we sample an in-context task $\mathcal{T}$ from dataset $\mathcal{D}$. Each task $\mathcal{T}$ consists of $K$ support examples $\mathcal{S}=\{(X_{s_{i}},Y_{s_{i}})\}_{i=1}^{K}$ and one query example $(X_{q},Y_{q})$, where $X_{q}$ and $X_{s_{i}}$ are images, and $Y_{q}$ and $Y_{s_{i}}$ are their one-hot semantic segmentation pseudo-labels. Among the $K$ support examples, one is a “positive” example sharing pseudo-classes with the query, while the remaining K1K-1 are “distractor” examples unlikely to share any pseudo-class. Since our method is unsupervised, these tasks and pseudo-labels are automatically constructed (see Sec. 3.2).
@@ 3.2 Automatic Dense In-Context Task Construction
Selecting examples for in-context tasks.  To generate in-context tasks, we pair each query image with one “positive” support image and KK-1 “distractor” support images.
To find positive images sharing similar visual content, we use DINOv2R global image representations. For each image, we retrieve five nearest neighbors using DINOv2R global features. This yields five positive pairs per image. We then retain pairs where the query and positive-support images share a common pseudo-class occupying more than $5\%$ of each image’s area. This ensures semantically meaningful positive pairs. The final refined list contains all possible query-positive support pairs for training.
For the KK-1 “distractor” examples, we randomly sample images from the current mini-batch during training. Including distractors encourages the model to distinguish relevant from irrelevant content, improving feature discriminability.
@@ 4.1 Experimental Setup
$\frac{1}{128}$$\frac{1}{64}$$\frac{1}{8}$$\frac{1}{1}$60606565707075758080Fraction of DatamIoU (%)(a) PascalVOC $\frac{1}{128}$$\frac{1}{64}$$\frac{1}{8}$$\frac{1}{1}$262628283030323234343636383840404242Fraction of Data(b) ADE20K
@@ 4.2 Retrieval-based Scene Understanding
Post-training other base models.  In Tab. 4, we evaluate in-context segmentation on ADE20K and PascalVOC using CLIP [61] and MAE [31] as base models, before and after DIP post-training. Our method consistently improves performance across all tested models, showing its versatility. MAE improves by +33.4+33.4 mIoU on PascalVOC. These gains transform MAE from weak to improved at in-context segmentation, highlighting our approach’s effectiveness in enhancing dense features.
@@ 5 Conclusion
We introduced DIP, a novel unsupervised post-training method that enhances dense image representations for in-context scene understanding. Leveraging meta-learning principles and automatically generating in-context tasks with pseudo-segmentations using Stable Diffusion, our approach avoids complex self-distillation architectures. DIP is computationally efficient (<<9h on an A100) and generalizes well across downstream dense retrieval tasks, including semantic segmentation and depth prediction. It outperforms both the initial pretrained vision encoder and prior state-of-the-art post-training methods, providing a simpler and more effective solution for improving dense representations in vision models.
@@ Appendix A Implementation details
During post-training, we fine-tune only the last three transformer blocks of the pretrained ViT encoder $f(\cdot)$ while keeping the remaining layers frozen. Our MLP projector $h(\cdot)$ consists of two linear layers with a non-linear activation function GELU [35]. Hidden feature dimension is set to $7D$ and fixed output dimension 6144. $\ell_{2}$-normalization is applied at the output of the MLP. We set the temperature parameter $\tau$ of the softmax operator to $0.07$ (see Tab. 7). Our ablation study demonstrates that the method is robust to the choice of temperature, with values ranging from $0.03$ to $0.09$ yielding similar performance on the PascalVOC dataset. During post-training our support set consists of 1 positive and 7 “distractor” examples (8 total examples). During training, we use the AdamW [49] optimizer with a learning rate of $2.25\times 10^{-7}$ and a weight decay of $0.05$. We train for $5$ epochs. We employ cosine learning rate schedule.
@@ Appendix B Additional quantitative results
$10^{5}$$10^{6}$$10^{7}$34343636383840404242Memory SizemIoU (%)(a) ViT-S/14 $10^{5}$$10^{6}$$10^{7}$34343636383840404242Memory Size(a) ViT-B/14
@@ 1 $\frac{1}{8}$ $\frac{1}{64}$ $\frac{1}{128}$
11 $\frac{1}{8}$ $\frac{1}{64}$ $\frac{1}{128}$ 11 $\frac{1}{8}$ $\frac{1}{64}$ $\frac{1}{128}$